# AudioCodecBench: A Comprehensive Benchmark for Audio Codec Evaluation

### v1

@@ Introduction
In the downstream probe tasks, to ensure fair and reliable results, all experiments are conducted under the same computational budget. For the ASR task, we select Speech2Text (Ott et al. 2019; Wang et al. 2020) as the tokenizer for word segmentation. The related tasks, datasets, and evaluation metrics are shown in Table 2. Detailed introductions are shown in Appendix DD.
The results of the music probe task are shown in Table 5. The visualized result is shown in Figure 22 in Appendix CC. In the ED task, SemantiCodec’s performance on Valence prediction is the best. This result supports the conclusion that Valence is closely associated with semantics (Asgari et al. 2014). Tasks such as MT, GC and KD involve high-level musical structures, SemantiCodec shows advantages in these tasks. Meanwhile, XCodec and SemantiCodec also achieved better performance in IC and PC tasks, which closely related to symbolic music information. Although tasks like SI and VTD rely on acoustic properties like timbre, codecs that combine both semantics and acoustics perform even better than acoustic codecs. This suggests that representation containing both semantics and acoustics makes information such as timbre more easily utilized by downstream models. Among these results, codecs that integrate semantics show better performance compared to acoustic codecs in these tasks. And these results also shows that semantics are crucial for the modeling of downstream tasks.
The results of the speech and sound probe tasks are shown in Table 6. The visualized result is shown in Figure 21 in Appendix CC. The SSL models achieve the best performance. Codecs that integrate semantics demonstrate the suboptimal performance. Acoustic codecs perform the worst. WavTokenizer achieves the lowest performance. In the ASR task, codecs that explicitly introduce semantics generally achieve better WER/CER scores than acoustic codecs. In the VSD task, codecs that combine both semantics and acoustics show outstanding performance. It further suggests that timbre information may be effectively retained and utilized in representations that contain both semantics and acoustics. In the ED task, the performance of different codecs is relatively balanced. This suggests that the emotion-related features required for this specific task can be fully fitted by codecs. Notably, in the ESC task, Mimi performs worse than other semantic codecs. This may be attributed to Mimi’s use of a larger hop length setting. This setting reduces its temporal resolution and weakens its ability to capture the transient features for ESC task.
Task Metric r
Reconstruction WERREC 0.06
@@
VSDACC 0.55
Table 7: Pearson correlation coefficient between PPL and metrics from various speech evaluation tasks.
In order to explore the impact of various metrics on the PPL value of LMs, we calculate the Pearson correlation coefficients between various task metrics and PPL. We aim to reveal which codec properties or audio features are more beneficial for LM modeling. The correlation coefficient is performed on the speech dataset and the results are shown in Table 7. PPL is positively correlated with CTC probe task metrics, which demonstrates that tokens rich in semantic content are easier for LMs to model. However, it shows a negative correlation with objective acoustic reconstruction metrics, indicating that overfitting acoustic details may increase the difficulty of LM modeling. Moreover, the reconstructed WER/CER metrics, which should be negatively correlated with PPL, show the opposite of our expected conclusion. We speculate that XCodec and YuE use different methods of integrating semantics. Their reconstruction metrics just below acoustic codecs like DAC and EnCodec. The codebook ID sensitivity metrics show a positive correlation with ppl, which indicates that the LM’s modeling capability is more dependent on the semantic content encoded in the token sequence.
@@ Appendix D Appendix D: Introduction to downstream probe tasks and related datasets
Instrument Classification (IC): This task aims to identify one or more musical instruments present in an audio recording. In the MARBLE classification system, this is considered an Acoustic-Level task, and its results evaluate the codec’s ability to represent fundamental acoustic properties like timbre. For evaluation, we use the NSynth dataset, which contains 11 single-instrument categories and is evaluated using Acc. We also use the MTG Instrument dataset, a multi-label collection with 41 labels, which is evaluated using ROC-AUC and PR-AUC/AP.

### v2

@@ Introduction
In the downstream probe tasks, to ensure fair results, all experiments are conducted under the same computational budget. For the ASR task, we select Speech2Text (Ott et al. 2019; Wang et al. 2020) as the text tokenizer. The related tasks, datasets, and evaluation metrics are shown in Table 2. Detailed introductions are shown in Appendix DD.
The results of the music probe task are shown in Table 5. The visualized result is shown in Figure 22 in Appendix CC. In the ED task, SemantiCodec’s performance on Valence prediction is the best. Arousal is more strongly associated with acoustic features, while Valence is more strongly associated with semantic content (Asgari et al. 2014). This is consistent with our results. Tasks such as MT, GC and KD involve high-level musical structures, SemantiCodec shows advantages in these tasks. Meanwhile, XCodec and SemantiCodec also achieved better performance in IC and PC tasks, which closely related to symbolic music information. These tasks share the common feature that their labels (e.g., “Pop,” “A major”) can be strictly described by text or symbols, with a correspondence between musical content and labels. We refer to these tasks as semantic-driven tasks. Therefore, in these tasks, semantic codecs show better performance than acoustic codecs. These results also validate our definition of “semantic,” proving that introducing semantics can effectively capture high-level, symbolizable information in music.
The results of the speech and sound probe tasks are shown in Table 6. The visualized result is shown in Figure 21 in Appendix CC. The SSL models achieve the best performance. Codecs that integrate semantics demonstrate the suboptimal performance. Acoustic codecs perform the worst. WavTokenizer achieves the lowest performance. In the ASR task, codecs that explicitly introduce semantics generally achieve better WER/CER scores than acoustic codecs. In the VSC task, codecs that integrate semantics show outstanding performance. It further suggests that timbre information may be effectively retained and utilized in representations that contain both semantics and acoustics. In the ED task, the performance of different codecs is relatively balanced. This suggests that the emotion-related features required for this specific task can be fully fitted by codecs.
Task Dataset Type Metric r
Reconstruction Speech WERREC 0.06
@@
VSCACC 0.55
ESCACC 0.67
EDACC 0.47
Music GCACC 0.2
[…]
EDAP(MTGMT) 0.43
EDRA(MTGMT) 0.58
MTAP(MTT) 0.59
MTRA(MTT) 0.49
MTAP(MTGT) 0.68
MTRA(MTGT) 0.73
ICACC(NS) 0.39
ICAP(MTGI) 0.65
ICRA(MTGI) 0.71
KDACC 0.62
VTDACC 0.41
PCACC 0.56
Table 7: Pearson correlation coefficient between PPL and metrics from various evaluation tasks.
In order to explore the impact of various metrics on LM modeling, we calculate the Pearson correlation coefficients between various task metrics and PPL. We aim to reveal which metrics or audio features are more beneficial for LM modeling. The correlation coefficient is performed on the results are shown in Table 7. PPL is positively correlated with CTC probe task metrics and very weak correlated with reconstructed WER/CER metrics, which demonstrates that tokens rich in semantic content are easier for LMs to model. However, it shows a negative correlation with objective acoustic reconstruction metrics, indicating that overfitting acoustic details may increase the difficulty of LM modeling. The ID sensitivity metrics show a positive correlation with PPL, which indicates that introducing semantics can bring more stable ID patterns, thereby benefiting the modeling of LMs.
@@ Appendix D Appendix D: Introduction to downstream probe tasks and related datasets
Instrument Classification (IC): This task aims to identify one or more musical instruments present in an audio recording. In the MARBLE classification system, this is considered an Acoustic-Level task, and its results evaluate the codec’s ability to represent fundamental acoustic features. For evaluation, we use the NSynth dataset, which contains 11 single-instrument categories and is evaluated using Acc. We also use the MTG Instrument dataset, a multi-label collection with 41 labels, which is evaluated using ROC-AUC and PR-AUC/AP.